Calculate the probability from the z-table. a. 𝑃(𝑍 > βˆ’0.45) b. 𝑃( 𝑍 < 2.10) c. 𝑃(0 < 𝑍 < 2.10) d. 𝑃(βˆ’0.45 < 𝑍 < 2.10)


1
Expert's answer
2021-06-24T09:08:26-0400

a. 𝑃(𝑍 > βˆ’0.45) = "P(-0.45<Z<0)+P(0<Z<\infty)=0.17364+0.5=0.67364"


b. 𝑃( 𝑍 < 2.10) = "P(-\infty<Z<0)+P(0<Z<2.10)=0.5+0.48214=0.98214"


c. 𝑃(0 < 𝑍 < 2.10) =0.48214


d. 𝑃(βˆ’0.45 < 𝑍 < 2.10)=


"P(-0.45<Z<0)+P(0<Z<2.10)=0.17364+0.48214=0.65578"